# Chain-of-Discussion: A Multi-Model Framework for Complex Evidence-Based Question Answering

### v1

@@ Abstract
Open-ended question answering requires models to find appropriate evidence to form well-reasoned, comprehensive and helpful answers. In practical applications, models also need to engage in extended discussions on potential scenarios closely relevant to the question. With augmentation of retrieval module, open-source Large Language Models (LLMs) can produce coherent answers often with different focuses, but are still sub-optimal in terms of reliable evidence selection and in-depth question analysis. In this paper, we propose a novel Chain-of-Discussion framework to leverage the synergy among multiple open-source LLMs aiming to provide more correct and more comprehensive answers for open-ended QA, although they are not strong enough individually. Our experiments show that discussions among multiple LLMs play a vital role in enhancing the quality of answers. We release our data and code at https://github.com/kobayashikanna01/Chain-of-Discussion.
@@ 1 Introduction
Figure 1: The process of Chain-of-Discussion, compared with chain-of-thought. The green parts are necessary to answer the user’s question. The blue parts indicate closely related to the question, which can be used for detailed discussion. The red parts are irrelevant content which should be avoided.
@@
Firstly, retrieval models are not entirely prefect, inevitably with noise in the retrieval results. Let us take legal consultation as an example. In Figure 1, the model is required to respond to a question regarding the necessity of child support payments. Due to the semantic similarity between obligations for supporting children (financially) and raising/protecting children (physically), the retrieval model may wrongly return law articles pertaining to guardianship qualifications. LLMs usually cannot filter all these noisy evidence, which may propagate and lead to incomplete analysis, wrong reasoning paths, biased opinions and finally problematic or even misleading answers.
@@
Secondly, we expect the LLMs output correct responses and consistent explanations, while providing more useful suggestions about the potential scenarios not directly mentioned in the questions but indeed helpful for users’ current or near-future situations. For instance, in Figure 1, when responding to a question about the obligation to pay child support for a user facing financial difficulties, the model should also remind her/him of the standards for child support payments and ways to negotiate for a reduction in the burden of child support given her/his current situation. This is even hard for humans where one should have access to proper evidence, e.g., the necessary or closely related law articles here, and accordingly provide kind reminders with reasonable explanations. Let alone LLMs without abundant annotations to train/fine-tune, which usually focus on the specific facts literally appearing in the questions.
@@
In this work, we will focus on the complex evidence-based question answering (CEBQA) task, a typical example of the open-ended QA tasks. We collect a high quality CEBQA dataset consisting of 200 carefully annotated legal consultation questions in the field of marriage and family affairs. To address the above challenges, we propose a novel chain-of-thought framework, the Chain-of-Discussion (CoD), which involves multiple LLMs in summarizing, criticizing and revising each other’s output to reach a well-supported and helpful response.
@@
Our motivations are two-fold. On the one hand, different LLMs may have different intrinsic knowledge and reasoning capabilities due to different training data. Thus, multiple LLMs can be less possible to make errors concurrently than a single LLM. Recent works Zhang et al. (2023) show checking the consistency across multiple LLMs helps reduce output hallucinations. Specifically, we propose a criticize-and-revise framework, which requires multiple LLMs to discuss and reach a consensus for a better response. For questions that need to involve helpful scenarios or possible extensions, we guess multiple LLMs may provide a diverse set of perspectives to address these possibilities. We thus propose a summarizing step to gather different but helpful perspectives from multiple LLMs, which will eventually form comprehensive and detailed responses based on the summarized analyses.
@@
Different from existing multi-model interaction works Chan et al. (2024); Zhang et al. (2023) using strong closed-source LLMs APIs, e.g., GPT-4 OpenAI (2023), we decide to take a challenge to study how to best exploit the small-scaled open-source LLMs, e.g., around 7B parameters, for a shared objective, while pushing the boundary of research regarding mult-model interaction.
@@ 2 Related Works
In this work, we face challenges more complex than RAG. While the model filters out irrelevant evidence, it also needs to retain evidence relevant to potential scenarios. Sometimes, determining which evidence can be used for potential scenarios and which is irrelevant evidence is also a challenging issue for humans.
Chain-of-Thought (CoT)
@@ 3 Preliminaries
In CEBQA tasks, given a user’s question $q$ and a store of evidence documents $\mathcal{D}$, a model should analyze $q$ first, find necessary evidence $\mathcal{D}_{q}=\{d_{1},\ \cdots,\ d_{t}\}$ from $\mathcal{D}$ accordingly and generate a paragraph $r$ as the final response. For instance, in the legal consultation task, users may ask what to do given her/his current situation. The model should find supportive evidence from a store of law articles, judicial interpretations, or previous legal cases, and generate a helpful and detailed response.
Specifically, we expect the generated responses to meet the requirements in terms of correctness and comprehensiveness. (1) Correctness: The responses should be based on the evidence that can support to answer the questions, and refrain from employing irrelevant evidence or misinterpreting the evidence out of context. (2) Comprehensiveness: The responses should engage in discussions about potential scenarios that would be relevant or helpful to the users, even if not explicitly mentioned in the users’ questions.
@@ 4 CoD: Summarize, Criticize, and Revise
Similar to the baseline, we employ a two-stage analyzing pipeline that instructs LLMs to analyze the question and evidence separately. To address the correctness and comprehensiveness of generated answers, at the stage of question analysis, we encourage models to read and summarize others’ analyses so as to take more scenarios closely relevant to the question into account, in the purpose of augmenting the comprehensiveness. During the stage of evidence analysis, we require all other LLMs to criticize the evidence analysis of each LLM. Subsequently, the model will read others’ critique and determine whether to revise its own analysis or not. The model finally generate a correct and more helpful response based on the summarized question analysis and revised evidence analysis.
@@ 4.1 Stage 1: Question Analysis
Different LLMs can have varying preferences in analyzing the potential scenarios. Therefore, we believe that by integrating the outputs of multiple LLMs, we can take more helpful scenarios into account, thus improve the comprehensiveness of question analysis. We prompt each LLM $M_{i}$ to analyze the question $q$, with retrieved evidence $\mathcal{D}^{q}$ as a reference: $a_{q,\,M_{i}}^{\text{que}}=f_{\text{que}}\left(q,\ \mathcal{D}_{q}|\theta_{M_{i}}\right).$
@@ 5 Experiments
As discussed in Section 3, legal consultation is a typical example for CEBQA tasks, which require model to generate an accurate response including helpful discussions about relevant scenarios.In our experiments, we delve into the legal consultation task in China, where all legal activities should be based on law articles and judicial interpretations, which can be naturally considered as the evidence store in our framework.
@@ 5.1 Data Collection
We focus on the legal consultation in the fields of marriage, family affairs, and inheritance, which cover various types of legal disputes such as divorce, custody, contracts, property and so on. We collect 200 questions from real users and the corresponding responses from consultants through Web Search Engines.
@@
To ensure the data quality, we manually check on the questions and answers. We correct all typos but retain the informal expressions in the questions. Note that there may be omissions or slight word-order inversions in the questions, which poses a challenge to the model’s reasoning capabilities.
We employ two annotators with background in civil law to examine the correctness and logical coherence of these responses. For the responses identified with errors, we encourage the two annotators to discuss and reach a consensus for modifications, otherwise, leave them as they are.
Evidence Annotation
We construct the evidence store based on all articles of the Civil Code and the Civil Procedure Law and their judicial interpretations. We categorize these article into three types: necessary, optional, and not required. The necessary articles are the ones highly relevant to the question, while the optional articles can be basis for the discussion of potential scenarios. Please see more details in Appendix B. We ensure there are 5 articles in each example. And on average, each example contains 1.52 necessary articles, 1.23 optional articles, and 2.25 not required articles. It means approximately 45% of the retrieved articles are not required at all.
@@ 5.2 Experimental Setup
In this work, we select open-source LLMs trained by different research groups. We hope these models have learnt different knowledge and gain different reasoning capabilities from pre-training. Then, these models may provide various analytical perspectives via interaction and compensate for deficiencies in their own reasoning capabilities.
We study four open-source fine-tuned LLMs, Baichuan2-7B Baichuan (2023), Deepseek-7B DeepSeek-AI (2024), Qwen-7B Bai et al. (2023), and Xverse-7B [^1] https://huggingface.co/xverse/XVERSE-7B-Chat, which are four of the best 7B-parameter LLMs performing on CMMLU Li et al. (2023a). When we use a specific LLM as the target model, the other three LLMs are expected to generate diverse question analyses and criticize the evidence analysis of target model.
@@
Single-model baselines (BS): Question analysis, article analysis, and response are all generated by a single LLM.
@@
Only Stage 2 (S2): Three other LLMs criticize the article analysis generated by target LLM. The question analysis and the final response are generated by target LLM on its own.
@@
We employ each LLM as the target model, replicating the experimental settings. We report the performance for each LLM as the target role. Please see more details in Appendix A.
@@
Different from the short-form open-domain QA whose answers are usually several words or sentences, the responses in the CEBQA tasks can consist of several hundred or even thousands of words. Therefore, it is impossible to employ the metrics such as F1 or exact match which are widely used in QA benchmarks Joshi et al. (2017); Rajpurkar et al. (2018). These generated responses are also complex which contain facts and causal relations to be verified. Thus, it is will be difficult to evaluate them all by humans due to unacceptable time costs.
Following previous works Liu et al. (2023); Chan et al. (2024), we employ GPT-4 to evaluate the quality of generated responses, with the human-written responses, necessary and optional articles as reference. We prompt gpt4-turbo-0125 to score the responses based on correctness and comprehensiveness. The scorer should assign integer scores between 1 and 10. If there is no clear reason to indicate that the responses of LLMs are significantly better or worse than human-written ones, a score of around 7 should be given. Please see the scoring prompts in Appendix E.
@@ 5.3 Main Results
Table 1: The average scores of each target LLM and each setting evaluated by GPT-4.
Table 1 shows the evaluation results produced by GPT-4. Comparing the results of baseline framework (BS) and Chain-of-Discussion (S1S2), we can find each LLM can obtain improvements from discussions with other LLMs, with Baichuan2-7B increased by +0.340, Deepseek-7B by +0.115, Qwen-7B by +0.120, Xverse-7B by + 0.110. We also find that employing multi-model discussion on both stages can bring more improvement than using it on one stage only.
We also have to acknowledge that although Chain-of-Discussion can enhance the LLMs, the CoD-augmented Baichuan2-7B, Qwen-7B, or Xverse-7B can still not outperform Deepseek-7B under baseline settings, with around 0.5 scores left behind. The results show that the quality of responses primarily relies on the inherent ability of LLM to comprehend contexts and then to generate.
We notice using multi-model discussion only at Stage 1 or Stage 2 fails to enhance Xverse-7B or Qwen-7B. We will provide more discussions and case studies in Section 6.3.
@@ 6.1 Evidence-Centric Evaluation
Different from previous Question Answering tasks Joshi et al. (2017); Kwiatkowski et al. (2019) whose answers can be several words or single sentences, in the CEBQA tasks, LLMs are required to provide both detailed and correct responses for the question and potential scenarios. Therefore, we wonder whether the Chain-of-Discussion framework can enhance the comprehensiveness and correctness of model output.
Similarly, when discussing the details of potential scenarios, LLMs should also reference optional evidence. Hence, we can assess the correctness and comprehensiveness of responses by the accuracy of reference to various types of evidence documents.
We propose two metrics of accuracy, N-Acc and O-Acc, to assess the correctness and comprehensiveness, respectively. We utilize the not required articles as negative samples. For N-Acc, we employ the necessary articles as positive samples, while the optional articles for O-Acc. We employ rule-based method to examine whether the response have used an article. Please see details in Appendix C.
@@
We compute the Macro average N-Acc and O-Acc across all examples. If an example does not contain optional articles, it will not participate in the calculation of O-Acc. Table 2 shows the results of each target LLM under different experimental settings. Compared to the baselines (BS), the Chain-of-Discussion framework (S1S2) can achieve around a 2% improvement on both N-Acc and O-Acc for Baichuan2-7B, Qwen-7B, and Xverse-7B. Even for Deepseek-7B, which performs the best in GPT-4-based evaluation, our proposed framework still brings improvements of 0.86% and 0.53% to N-Acc and O-Acc, respectively. Recalling the GPT-4-evaluated results in Table 1, where Baichuan2-7B obtains the most improvement of overall quality. We find this LLM also get the most improvement on N-Acc and O-Acc, with improvements of 4.91% and 2.24%, respectively.
@@
The results indicate that introducing multi-model discussions during both question analysis and evidence analysis contributes to increasing the probability of LLMs referencing correct evidence. It can be one of the reasons why Chain-of-Discussion can improve the quality of model responses.
@@
Comparing the results under the setting of BS, S1, and S2, we can find that involving multiple LLMs in a single stage can actually enhance both correctness and comprehensiveness. However, overall, employing multi-model discussions in question analysis contributes more to the comprehensiveness, while introducing other models in evidence analysis brings more improvement in correctness.
Question: What is the difference between resumption of marital relationship (复婚) and remarriage with other person (再婚)?
6.2 Manual Check
To further study the quality of responses generated by CoD, we randomly sample 30 cases and manually examine the responses in terms of fluency and logicality. We select the responses generated by Qwen-7B or Qwen7B+CoD, which get worse average GPT-4-evaluated scores than responses of other LLMs. We find Qwen-7B with vanilla CoT has a poor ability to comprehend the articles, often resulting in logical errors in responses, while the CoD mechanism introduces opinions and critics from other LLMs, thus helps to distinguish ambiguous terms, and reach better logicality.
Table 3 show the case where CoD can help to reduce logical errors in Qwen-7B’s responses. We find Qwen-7B fails to understand that both 复婚 (reconcile and remarry) and 再婚 (remarry with other person) are considered as marriage in legal terms, which should comply with the provisions of Article 1046 but not Article 1083. We believe that discussions with other models can, to some extent, reduce the hallucination caused by the target model’s poor reasoning capabilities.

### v2

@@ Abstract
Open-ended question answering requires models to find appropriate evidence to form well-reasoned, comprehensive and helpful answers. In practical applications, models also need to engage in extended discussions on potential scenarios closely relevant to the question. With augmentation of retrieval module, open-source Large Language Models (LLMs) can produce coherent answers often with different focuses, but are still sub-optimal in terms of reliable evidence selection and in-depth question analysis. In this paper, we propose a novel Chain-of-Discussion framework to leverage the synergy among multiple open-source LLMs aiming to provide more correct and more comprehensive answers for open-ended QA, although they are not strong enough individually. Our experiments show that discussions among multiple LLMs play a vital role in enhancing the quality of answers. We will release our data and code for further research.
@@ 1 Introduction
Figure 1: The process of Chain-of-Discussion, compared with chain-of-thought. The green parts are necessary to answer user’s question. Blue parts indicate closely related to the question, useful for detailed/extended discussions. Red parts are irrelevant content that should be avoided.
@@
Firstly, retrieval models are not entirely prefect, inevitably with noise in the retrieval results. As a legal consultation example shown in Figure 1, the model is required to respond to a question regarding the necessity of child support payments. Due to the semantic similarity between obligations for supporting children (financially) and raising/protecting children (physically), the retrieval model may wrongly return law articles pertaining to guardianship qualifications. LLMs usually cannot filter all these noisy evidence, which may propagate and lead to incomplete analysis, wrong reasoning paths, biased opinions and finally problematic or even misleading answers.
@@
Secondly, we expect LLMs output correct responses and consistent explanations, providing useful suggestions about potential scenarios not directly mentioned in questions but indeed helpful for users’ current/near-future situations. In Figure 1, when responding to a question about the obligation to pay child support for a user facing financial difficulties, the model should also remind her/him of the standards for child support payments and ways to negotiate for a reduction given her/his current situation. This is even hard for humans where one should have access to proper evidence, e.g., necessary or closely related law articles, and accordingly provide kind reminders with reasonable explanations. Let alone LLMs without specific training/fine-tuning, which usually focus on the specific facts literally appearing in the questions.
@@
In this work, we focus on the complex evidence-based question answering (CEBQA) task, a typical example of the open-ended QA tasks. We collect a high quality CEBQA dataset consisting of 200 carefully annotated legal consultation questions in the field of marriage and family affairs. To address the challenges, we propose a novel chain-of-thought (CoT) framework, Chain-of-Discussion (CoD), encouraging multiple LLMs to summarize, criticize and revise each other’s output to reach a well-supported and helpful response.
@@
Our motivations are two-fold. First, different LLMs may have different intrinsic knowledge and reasoning capabilities due to different training data. Thus, multiple LLMs can be less possible to make errors concurrently than a single LLM. Recent works Zhang et al. (2023) show checking the consistency across multiple LLMs helps reduce output hallucinations. Specifically, we propose a criticize-and-revise framework, which requires multiple LLMs to discuss and reach a consensus for a better response. For questions that need to involve helpful scenarios or possible extensions, we guess multiple LLMs may provide a diverse set of perspectives to address these possibilities. We thus propose a summarizing step to gather different but helpful perspectives from multiple LLMs, which will eventually form comprehensive and detailed responses based on the summarized analyses.
@@
Different from existing multi-model interaction works Chan et al. (2024); Zhang et al. (2023) using strong closed-source LLMs, e.g., GPT-4 OpenAI (2023), we decide to take a challenge to study how to best exploit the small-scaled open-source LLMs, e.g., around 7B parameters, for a shared objective, while pushing the boundary of research regarding multi-model interaction.
@@ 2 Related Works
In this work, we face more complex challenges than RAG. While the model filters out irrelevant evidence, it also needs to retain evidence relevant to potential scenarios. Sometimes, determining which evidence can be used for potential scenarios and which are irrelevant is also challenging for humans.
Chain-of-Thought
@@ 3 Preliminaries
In CEBQA tasks, given a user’s question $q$ and a store of evidence documents $\mathcal{D}$, a model should analyze $q$ first, find necessary evidence $\mathcal{D}_{q}=\{d_{1},\ \cdots,\ d_{t}\}$ from $\mathcal{D}$ accordingly and generate a paragraph $r$ as the final response. For instance, in the legal consultation task, users may ask what to do given her/his current situation. The model should find supportive evidence from a store of law articles or previous legal cases, and generate a helpful and detailed response.
Specifically, we expect the generated responses to meet the requirements in terms of correctness and comprehensiveness. (1) Correctness: The responses should be based on the evidence that can support to answer the questions, and refrain from employing irrelevant evidence or misinterpreting evidence out of context. (2) Comprehensiveness: The responses should engage in discussions about potential scenarios that would be relevant/helpful to users, even not explicitly mentioned in questions.
@@ 4 CoD: Summarize, Criticize, and Revise
Similar to the baseline, we employ a two-stage analyzing pipeline that instructs LLMs to analyze the question and evidence separately. To address the correctness and comprehensiveness of generated answers, during question analysis, we encourage models to read and summarize others’ analyses so as to consider more scenarios closely relevant to the question, in the purpose of augmenting the comprehensiveness. During the stage of evidence analysis, we require all other LLMs to criticize the evidence analysis of each LLM. Subsequently, the model will read others’ critique and determine whether to revise its own analysis or not. The model finally generate a correct and more helpful response based on the summarized question analysis and revised evidence analysis.
@@ 4.1 Stage 1: Question Analysis
Different LLMs can have varying preferences in analyzing the potential scenarios. Therefore, we believe that by integrating the outputs of multiple LLMs, we can take more helpful scenarios into account, thus improve the comprehensiveness of question analysis. We prompt each LLM $M_{i}$ to analyze the question $q$, with retrieved evidence $\mathcal{D}_{q}$ as a reference: $a_{q,\,M_{i}}^{\text{que}}=f_{\text{que}}\left(q,\ \mathcal{D}_{q}|\theta_{M_{i}}\right).$
@@ 5 Experiments
As there is no existing dataset for CEBQA tasks, here we delve into the legal consultation task and collect the first CEBQA dataset. In civil law systems like China, all legal activities, including legal consultation, should be based on law articles (or judicial interpretations), which can be naturally considered as the evidence store in our framework.
@@ 5.1 Data Collection
To reflect the diversity in practical scenarios, we focus on the fields of marriage, family affairs, and inheritance, covering various legal disputes, e.g., divorce, custody, contracts, property, etc. From a pool of 609 questions collected from real users with consultants’ responses through Web Search Engines, we employ an legal expert to select 200 questions to ensure their semantic distinctiveness and coverage. See more details in Ethics Statement.
Evidence Annotation
We construct the evidence store based on all 1,749 articles of the Civil Code, Civil Procedure Law and their judicial interpretations in China. For each question, we consider three types of articles: necessary, optional, and not required. Necessary articles are highly relevant to the question, while optional ones can be basis for discussions of potential scenarios (more details in Appendix B). We ensure there are 5 articles retrieved for each question, and on average, each contains 1.52 necessary, 1.23 optional, and 2.25 not required articles. It means about 45% of the retrieved articles are not required at all.
@@
We employ 6 annotators with background in civil law to manually check the questions, answers and articles. Annotators are instructed to correct all typos but retain the informal expressions in questions. Note that there are many omissions or slight word-order inversions in the questions, posing a challenge to models’ reasoning capabilities.
Annotators also examine the correctness and logical coherence of the responses. For problematic ones, annotators are encouraged to discuss and reach a consensus for modifications, otherwise, leave them as they are. Averagely, it takes about 20 hours per annotator to examine 100 instances.
@@ 5.2 Experimental Setup
As our main focus is to investigate whether small open-source LLMs can collaborate through summarize-criticize-revise, we study four open-source fine-tuned LLMs, Baichuan2-7B Baichuan (2023), Deepseek-7B DeepSeek-AI (2024), Qwen-7B Bai et al. (2023), and Xverse-7B [^1] https://huggingface.co/xverse/XVERSE-7B-Chat, which are four of the best 7B-parameter LLMs performing on CMMLU Li et al. (2023a). When we use a specific LLM as the target model, the other three LLMs are expected to generate diverse question analyses and criticize the evidence analysis of target model.
To examine if close-source LLMs can still benefit from our CoD, we also test with gpt-3.5-turbo-1106, gemini-1.0-pro-latest, and claude-3-haiku-20240307 similarly to the open-source group.
@@
Single-model baselines (BS): One LLM takes a query and several retrieved articles as input and performs question analysis, article analysis, and response generation in a vanilla CoT manner.
@@
Only Stage 2 (S2): Three other LLMs criticize the article analysis generated by the target LLM. The question analysis and the final response are generated by the target LLM on its own.
@@
We employ each LLM as the target model, replicating the experimental settings. We report the performance for each LLM as the target role. More details and prompt templates are in Appendix A and  E.
@@
Different from conventional QA tasks, the responses in CEBQA tasks can consist of several hundred or even thousands of words, which are also knowledge intensive and complex in structures, containing facts and causal relations to be verified. Therefore, it is impossible to employ the popular metrics such as F1 or exact match Joshi et al. (2017); Rajpurkar et al. (2018).
Following previous works Liu et al. (2023); Chan et al. (2024), we employ GPT-4 to evaluate the quality of generated responses, with the expert-written responses, necessary and optional articles as reference. We prompt gpt4-turbo-0125 to score a response in an integer between 1 and 10 based on correctness and comprehensiveness. If there is no clear reason to indicate a response is significantly better or worse than human-written ones, a score of around 7 should be given (Scoring prompt is in Appendix F).
@@ 5.3 Main Results
Close-source LLMs Group
GPT-3.5-turbo BS 6.895
S1S2 6.955 +0.060
Gemini-1.0-pro BS 6.940
S1S2 6.975 +0.035
Claude-3-haiku BS 7.280
S1S2 7.300 +0.020
Table 1: The average scores of each target LLM and each setting evaluated by GPT-4. The upper is evaluated within the open-source LLMs group, while the bottom is within the closed-source LLMs group.
Table 1 shows the evaluation results produced by GPT-4. Comparing the baseline CoT setting (BS) and Chain-of-Discussion (S1S2) in the open-source group, we can find each LLM can obtain improvements from discussions with other LLMs, with Baichuan2-7B increased by +0.340, Deepseek-7B by +0.115, Qwen-7B by +0.120, Xverse-7B by + 0.110. We think other LLMs bring more related aspects according to their own strengths into discussions while the baseline COT setting has to rely on one LLM only. We also find employing multi-model discussion on both stages can bring more improvement than using it on one stage only.
Although CoD can enhance all LLMs, the CoD-augmented Baichuan2-7B, Qwen-7B, or Xverse-7B still can not outperform Deepseek-7B under its baseline setting, with around 0.5 left behind. This may indicate that the quality of responses primarily relies on the inherent ability of LLMs to analyze context and then to generate.
As shown in the bottom of Table 1, even for powerful closed-source LLMs like Claude-3, our CoD framework can still brings improvement consistently, but the improvements are admittedly small compared to the open-source group.
Through a manual check, we also find that these powerful closed-source LLMs may also produce erroneous analyses of questions, resulting in incorrect responses. We believe that regardless the model scales, our CoD framework can leverage the varied knowledge and capabilities of multiple models to achieve better performance.
@@ 6.1 Evidence-Centric Evaluation
Besides overall evaluation by GPT-4, we wonder if our Chain-of-Discussion framework can enhance the comprehensiveness and correctness of the model output. When discussing the details of questions or potential scenarios, LLMs should refer to necessary or optional evidence. Hence, we can assess the correctness and comprehensiveness of responses by the accuracy of evidence documents.
We accordingly design two metrics of accuracy, N-Acc and O-Acc, to assess the correctness and comprehensiveness, respectively. We utilize the not required articles as negative samples. For N-Acc, we employ the necessary articles as positive samples, while the optional articles for O-Acc. We employ heuristic methods to examine if a response has used an article (details in Appendix C).
@@
We compute the Macro average N-Acc and O-Acc across all examples, shown in Table 2. We can see these LLMs exhibit differently on their own, and when performing a simple majority vote among all LLMs, we could get 77.62% for N-Acc, confirming our hypothesis that it is possible to harness the synergy among multiple open-source LLMs for a better performance. Compared to the baselines (BS), our CoD (S1S2) can achieve around a 2% improvement on both N-Acc and O-Acc for Baichuan2-7B, Qwen-7B, and Xverse-7B. Even for Deepseek-7B, the best in GPT-4 based evaluation, our framework still brings improvements of 0.86% and 0.53% to N-Acc and O-Acc, respectively.
@@
The results indicate that introducing multi-model discussions during both question analysis and evidence analysis helps target LLMs better refer to correct evidence. This also explains why CoD can improve the quality of model responses.
@@
Comparing the results under BS, S1, and S2, we find that involving multiple LLMs in a single stage can actually enhance both correctness and comprehensiveness. Specifically, employing multi-model discussions in question analysis contributes more to comprehensiveness, while introducing other models in evidence analysis (BS vs. S2) helps filtering irrelevant evidence thus brings more improvement in correctness.
6.2 Human Evaluation
We further examine the quality of generated responses through a win-rate analysis by human experts. We randomly sample 50 examples, and ask two human experts to examine whether the responses of CoD-enhanced models are better than those of baselines. We use the responses of Baichuan2-7B and Xverse-7B, which are models obtaining the most and the least improvement from CoD, respectively.
Figure 2: Human preference evaluation, comparing the CoD settings of Baichuan2-7B and Xverse-7B to their corresponding baseline settings across 50 randomly sampled examples.
As shown in Figure 2, in 38% of the examples, CoD can bring improvements to the responses of Baichuan2-7B, while 40% responses generated by Xverse+CoD surpass Xverse’s baseline setting. For Baichuan2-7B, only 16% of responses deteriorated after introducing the CoD framework, while for Xverse, only 12% worsened.
Qualitative Analysis. We also manually examine 30 responses of both models. The main error types are misunderstanding key legal terms (60%), discussing not-required articles (15%), and inconsistent explanations (10%). The latter is often caused by term misunderstanding. This shows that there may not be sufficient legal knowledge in 7B models to distinguish very similar terms. Comparing our CoD and baseline setting, we can see CoD can correct improper articles for 6 out 30 cases, leading to better results, through the cooperation mechanism.
Correlation between GPT-4 Evaluation and Human Although GPT-4 has been deemed as a reliable evaluator in various LLM evaluations Liu et al. (2023), we still want to examine how well GPT-4 preferences correlate to human experts in our challenging task. For each example from the randomly sampled 50 cases: (1) if GPT-4 assigns a higher score for the response generated by the CoD setting, we denote this evaluation result as +1; (2) if GPT-4 assigns a higher score for the response by the baseline setting, we denote this evaluation result as -1; Otherwise, we denote it as 0. We use the same principle to annotate the evaluation results of human expert.
The Pearson correlation coefficient between GPT-4 scores and human expert evaluation is 0.5863 with $p=7.7\times 10^{-6}$. This indicates a relatively high correlation between the two evaluation methods, where we could consider GPT-4 evaluation as a reliable manner for our work.